OpenStudy (anonymous):

I don't understand the steps ; Please help @amistre64 Use the general solution to solve 5 – 6x = 8x + 17

OpenStudy (xxcoyotekidxx):

you just got answers @Miracrown and i can help better than mira can if you will take my advice

OpenStudy (xxcoyotekidxx):

I am in Algebra II as a sophomore in high school after all

OpenStudy (anonymous):

I'm in Algebra 1 ; 9th grade

OpenStudy (xxcoyotekidxx):

i was in it in 8th grade if you'll let me help I can show all the steps so as to explain it

OpenStudy (anonymous):

Ok. Please .. i don't understand at all

OpenStudy (anonymous):

is -26 one of your answers or no ?

OpenStudy (nirmalnema):

5 – 6x = 8x + 17 5-17=8x+6x -12=14x x=-12/14 =-6/7

OpenStudy (anonymous):

there isn't any choices.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@nirmalnema How did you do that ?

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

@nirmalnema You were correct . :)

OpenStudy (xxcoyotekidxx):

first you subtract 17 from both sides -12-6x=8x then add 6x to both sides -12=14x then you divide both sides by 14 -12/14=x and that simplifies to -6/7=x

OpenStudy (nirmalnema):

in 1st step add 6x and subtract 17 both sides

OpenStudy (anonymous):

Thank you soo much everyone who helped! @nirmalnema Can you give @XxCoyoteKidxX a medal

OpenStudy (xxcoyotekidxx):

@nirmalnema you have to show all of them b/c some teachers count that as 2 steps

OpenStudy (nirmalnema):

ohh ok.. yaa i have given @justsmile531

OpenStudy (xxcoyotekidxx):

thanks guys